The cost of retained earnings the required rate of If a firm cannot invest retained earnings to earn a rate of return returnThe cost of equity using the discounted cash flow (or dividend-yield-plus-growth-rate) approach Tyler Enterprisess stock is
Answer #1

a) higher than

b) =3%+1.13*8.17% =12.23%

c) =9.88%+2.08% =11.96%

d) =2.75/20+6.4% =20.15%

e) =60%*15.5% =9.3%